OpenStudy (anonymous):

At maximum speed an airplane travels 1720 miles against the wind in 5 hours. Flying with the wind, the plane can travel the same distance in 4 hours. Let X be the maximum speed of the plane and Y be the speed of the wind. What is the speed of the plane with no wind?

OpenStudy (anonymous):

@myko, I have four questions that I need the answers to.

OpenStudy (anonymous):

i think @soati is giving you the answer, :)

OpenStudy (anonymous):

(X-Y) 5 =1720 That is, the speed of the plane minus the speed of the wind over 5 hours gives you that distance (X+Y) 4 = 1720 Similar reasoning. (X-Y)5 = 1720 (X+Y)4 = 1720 Lets solve the first equation first. You get X - Y = 344 so X = 344 + Y So we take that and replace X on the second equation by 344 + Y (344 + Y + Y)4 = 1720 344 + 2Y = 430 2 Y = 86 Y = 43. Thats the speed of the wind. Now we plug that value of Y on the equation we got for X earlier, resulting in X = 344 + 43 = 387. So that, 387, is the velocity of the plane without the effect of the wind.

OpenStudy (anonymous):

1720/x=5 5x=1720 x=1720/5 x=344 mph against the wind. 1720/y=4 4y=1720 y=1720/4 y=430 mph. with the wind. (430-344)/2 86/2=43 mph. is the speed of the wind. 344+43=387 430-43=387 mph. speed of the plane in still air.

OpenStudy (anonymous):

damn im too late for the partay